OpenStudy (anonymous):

Find the radius of convergence of the Taylor series around x=0 for \[\ln (1/(1-7x))\]

OpenStudy (anonymous):

can you show some steps,please?

OpenStudy (anonymous):

\[\frac 1 {1- 7x}= \sum_{n=1}^\infty (7x)^n \] \[\ln(1- 7x)=\int_0^x \frac {dt }{1- 7t}= \sum_{n=1}^\infty \int_0^x (7t)^n dt= \sum_{n=1}^\infty \frac { (7x)^{n +1}} {n+1}dt \]

OpenStudy (anonymous):

\[ \ln \frac 1 { 1- 7x}= \ln 1 - \ln (1- 7x)= -\ln (1-7x) \] From th post above let \[ c_n = \frac {7^{n+1}} {n+1} \] \[\lim_{n\to \infty} \frac {c_n}{c_{n+1}}= \frac 1 7 \]

OpenStudy (anonymous):

how do we know the first step?

OpenStudy (experimentx):

i guess i understand now

OpenStudy (anonymous):

can you kind of explain?

OpenStudy (experimentx):

which part?

OpenStudy (anonymous):

the first step

OpenStudy (experimentx):

1/(1-7x) = summation (7x)^n

OpenStudy (experimentx):

this is done using binomial expansion

OpenStudy (anonymous):

oh,we didn't cover that!!

OpenStudy (experimentx):

(1-7x)^-1 = 1+(-1)/1!*(-7x)+(-1)(-2)/2!(-7x)^2+(-1)(-2)(-3)/3!(-7x)^3+.. i guess it can be done using Tylor series too

OpenStudy (anonymous):

yea,i think i got it now!!thanks!!